OpenStudy (anonymous):

Ms. Carr's class is selling magazines to raise money for a field trip. The students in the class want to decide how much profit they'll make from each magazine they sell. The following formula describes this scenario: \[p=\frac{ m }{ n }\] p = Profit made from each magazine sold m = Money they want to raise n = Number of magazines sold Rewrite the formula to determine the number of magazines they'll need to sell. @texaschic101

OpenStudy (anonymous):

n=m/p?

OpenStudy (texaschic101):

I believe you are looking for n p = m/n (multiply the equation by n) np = m (now divide by p) n = m/p I agree with Kristen :)

OpenStudy (anonymous):

Thanks @Kristen17 and @texaschic101 you were both right! :)

OpenStudy (texaschic101):

happy to hear :)

OpenStudy (anonymous):

No problem:)